OpenStudy (anonymous):

Without integrating, state the integration formula you can use to integrate each of the following. ∫(e^x)/(e^x+1) ∫xe^x^2 dx

myininaya (myininaya):

use u substitution for both of them 1st one let u=e^x+1 and the second let u=x^2